OpenStudy (anonymous):

How do you find the max value, min value, and value of theta at max if the equation is y= 4sin (1/2θ-π/3) without graphing?

OpenStudy (anonymous):

Well, you know that the maximum value of sin(theta)=1 so the max value would occur when sin(1/2θ-π/3)=1. It is a similar concept for the min value. The min value of sin(theta) is 0 so the function is minimized when sin is 0. These concepts work because the function is just a multiplication.

OpenStudy (anonymous):

For the third part, you find the value of theta that makes sin max or 1. That is when (1/2θ-π/3)=pi/2 b/c sin is one at pi/2

OpenStudy (anonymous):

Isn't the minimum value of sin theta -1? And thank you so much.

OpenStudy (anonymous):

oh right. Yeah